# The inverse initial data problem for anisotropic Navier–Stokes equations via Legendre time reduction method

### v2

@@ 3 The time-dimensional reduction method
For large $N$, due to Theorem 3.5, the spectral projection $\mathbb{P}^{N}$ approximatedly commutes with $\partial_{t}$ and with $({\bf u}\cdot\nabla){\bf u}$. In other words, “project-then-operate” and “operate-then-project” agree in the limit. This consideration motivates our choice of the exponentially weighted Legendre basis $\{\Psi_{n}\}_{n\geq 0}$. Other time bases (e.g., the exponentially weighted polynomial basis in $L^{2}(0,T)$; see [13, 28]) need not have this property and may therefore be unsuitable for deriving (3.12).
@@
$\mathbb{P}^{N}[\mathbf{u}]({\bf x},t)=\sum_{n=0}^{N}\mathbf{u}_{n}({\bf x})\Psi_{n}(t).$ (3.13)
Substituting (3.13) and into the system (3.11), we obtain the following approximation:
$\sum_{n=0}^{N}\mathbf{u}_{n}({\bf x})\Psi_{n}^{\prime}(t)+\sum_{l=0}^{N}\sum_{n=0}^{N}\left(\mathbf{u}_{l}({\bf x})\cdot\nabla\right)\mathbf{u}_{n}({\bf x})\Psi_{l}(t)\Psi_{n}(t)=-\nabla p+\sum_{n=0}^{N}{\rm div}\left(\bm{\mu}:\nabla\mathbf{u}_{n}({\bf x})\right)\Psi_{n}(t),$ (3.14)
valid for all $({\bf x},t)\in\Omega\times(0,T)$. To derive the time-reduced model, we test equation (3.14) against $e^{-2t}\Psi_{m}(t)$ and integrate over $t\in(0,T)$ for each $m=0,1,\dots,N$, resulting in:
$\sum_{n=0}^{N}s_{mn}\mathbf{u}_{n}({\bf x})+\sum_{l=0}^{N}\sum_{n=0}^{N}a_{mnl}\left(\mathbf{u}_{l}({\bf x})\cdot\nabla\right)\mathbf{u}_{n}({\bf x})=-\nabla p_{m}({\bf x})+{\rm div}\left(\bm{\mu}:\nabla\mathbf{u}_{m}({\bf x})\right),$ (3.15)
for all ${\bf x}\in\Omega$, where the coupling coefficients are defined by
$s_{mn}=\int_{0}^{T}e^{-2t}\Psi_{n}^{\prime}(t)\Psi_{m}(t)\,dt,\quad a_{mnl}=\int_{0}^{T}e^{-2t}\Psi_{l}(t)\Psi_{n}(t)\Psi_{m}(t)\,dt,$
@@
$p_{m}({\bf x})=\int_{0}^{T}e^{-2t}p({\bf x},t)\Psi_{m}(t)dt.$
The corresponding Cauchy boundary conditions for each mode $\mathbf{u}_{m}$ follow from (1.2) and (3.12):
$\mathbf{u}_{m}({\bf x})=\textbf{0}\quad\partial_{\nu}\mathbf{u}_{m}({\bf x})=\int_{0}^{T}e^{-2t}\mathbf{f}({\bf x},t)\Psi_{m}(t)\,dt,$ (3.16)
for all ${\bf x}\in\partial\Omega$. Equations (3.15) and (3.16) define a coupled boundary value problem for the vector of coefficients $\mathbb{F}^{N}[\mathbf{u}]$. We refer to this as the time-dimensional reduction model:
$\begin{cases}\displaystyle\sum_{n=0}^{N}s_{mn}\mathbf{u}_{n}({\bf x})+\sum_{l=0}^{N}\sum_{n=0}^{N}a_{mnl}\left(\mathbf{u}_{l}({\bf x})\cdot\nabla\right)\mathbf{u}_{n}({\bf x})=-\nabla p_{m}+{\rm div}\left(\bm{\mu}:\nabla\mathbf{u}_{m}({\bf x})\right),&{\bf x}\in\Omega,\\ \displaystyle\mathbf{u}_{m}({\bf x})=\textbf{0},&{\bf x}\in\partial\Omega,\\ \displaystyle\partial_{\nu}\mathbf{u}_{m}({\bf x})=\mathbf{f}_{m}({\bf x}):=\int_{0}^{T}e^{-2t}\mathbf{f}({\bf x},t)\Psi_{m}(t)\,dt,&{\bf x}\in\partial\Omega,\end{cases}$ (3.17)
@@
Solving the system (3.17) yields the vector of Fourier modes $\mathbb{F}^{N}[\mathbf{u}]$, which can be used to reconstruct an approximation of the original space-time solution $\mathbf{u}({\bf x},t)$ via the expansion $\mathbb{P}^{N}[\mathbf{u}]=\mathbb{S}^{N}[\mathbb{F}^{N}[\mathbf{u}]]$. According to the approximation result in (3.13), this reconstruction provides a full space-time representation of $\mathbf{u}$, thereby transforming the inverse problem under consideration (Problem 1) into the computational task of solving the reduced model (3.17). Despite this reduction, solving the system (3.17) remains highly nontrivial due to the presence of the nonlinear convective term
$\sum_{l=0}^{N}\sum_{n=0}^{N}a_{mnl}\left(\mathbf{u}_{l}({\bf x})\cdot\nabla\right)\mathbf{u}_{n}({\bf x}),\quad m\in\{0,1,\dots,N\},$ (3.18)
making (3.17) a coupled nonlinear system. We will apply a conventional Picard iteration to overcome this difficulty, see Section 4.
@@
The time-dimensional reduction model (3.17) is an approximate representation of the original system, stemming from the truncation of the infinite Fourier expansion in (3.13). Despite this, we adopt the working assumption that
$\mathbf{U}=\begin{bmatrix}\mathbf{u}_{0}&\mathbf{u}_{1}&\dots&\mathbf{u}_{N}\end{bmatrix}^{\top}$
satisfies the reduced model. This assumption is justified by the fact that the truncation error diminishes as the cutoff parameter $N$ increases, with the error terms in the left-hand sides of (3.4) and (3.5) vanishing in the limit as $N\to\infty$, see Theorem 3.5. Consequently, the reduced model becomes an increasingly accurate surrogate for the full dynamics. Moreover, from a computational standpoint, the assumption is natural: practical implementations inherently involve a finite $N$, making the exact infinite-dimensional formulation intractable.
@@
The uniqueness of the time-reduction model (3.17) remains an open problem due to the complexity introduced by the anisotropic fourth-order tensor $\bm{\mu}$ and the presence of the nonlinear convection term given in (3.18). A well-established approach in the literature for establishing uniqueness in inverse problems is the use of Carleman estimates. However, deriving a Carleman estimate in the fully anisotropic setting is highly nontrivial and currently lacks a general theoretical framework. In contrast, if the problem is restricted to the isotropic case, Carleman estimates become more tractable; see, for instance, [9, 11]. Under such simplifications, the uniqueness of the time-reduction model (3.17) could then be rigorously analyzed. Therefore, we reserve a detailed investigation of uniqueness and stability results based on Carleman estimates for the isotropic case as a direction for future research.
4 A Picard algorithm to solve the time-reduction model
@@
which contains the solution of the reduced system in (3.17). Let the initial guess of the true solution be denoted by
@@
is known for some $k\geq 0$, we define the next iterate
${\bf U}^{(k+1)}=\begin{bmatrix}{\bf u}^{(k+1)}_{0}&{\bf u}^{(k+1)}_{1}&\dots&{\bf u}^{(k+1)}_{N}\end{bmatrix}^{\top}$
as the solution to the following boundary value problem:
@@
$\begin{cases}\displaystyle\sum_{n=0}^{N}s_{mn}\mathbf{u}_{n}^{(k+1)}({\bf x})+\sum_{l=0}^{N}\sum_{n=0}^{N}a_{mnl}\left(\mathbf{u}_{l}^{(k)}({\bf x})\cdot\nabla\right)\mathbf{u}_{n}^{(k)}({\bf x})\\ \hskip 142.26378pt=-\nabla p_{m}+{\rm div}\left(\bm{\mu}:\nabla\mathbf{u}_{m}^{(k+1)}({\bf x})\right),&{\bf x}\in\Omega,\\ \mathbf{u}_{m}^{(k+1)}({\bf x})=\mathbf{0},&{\bf x}\in\partial\Omega,\\ \partial_{\nu}\mathbf{u}_{m}^{(k+1)}({\bf x})=\mathbf{f}_{m}({\bf x}),&{\bf x}\in\partial\Omega,\end{cases}$ (4.1)
for each $m=0,1,\dots,N$. This iteration scheme is inspired by the classical Picard method. In general, the convergence of the sequence $\{{\bf U}^{(k)}\}_{k\geq 0}$ depends on the proximity of the initial guess ${\bf U}^{(0)}$ to the true solution of (3.17). However, we do not pursue this local convergence analysis, as such accurate initial guesses are typically unavailable in practice. Instead, we demonstrate numerically that the iteration converges even when the initial guess is chosen as ${\bf U}^{(0)}=\mathbf{0}$.
@@
The next step is to solve the over-determined problem (4.1). By over-determined, we mean the presence of both Dirichlet and Neumann information of ${\bf U}^{(k+1)}$ in this problem. We propose to use the quasi-reversibility method, first introduced in [18]. That means the solution ${\bf U}^{(k+1)}$ is the minimizer of a least-squares functional involving a Tikhonov regularization term
@@
$J^{\epsilon,N}_{{\bf U}^{(k)}}(\bm{\varphi})=\sum_{m=0}^{N}\Bigg[\int_{\Omega}\Big|{\rm div}\left(\bm{\mu}:\nabla\varphi_{m}({\bf x})\right)-\sum_{n=0}^{N}s_{mn}\varphi_{n}({\bf x})-\sum_{l=0}^{N}\sum_{n=0}^{N}a_{mnl}\left({\bf u}_{l}^{(k)}({\bf x})\cdot\nabla\right)\mathbf{v}_{n}^{(k)}({\bf x})\\ -\nabla p_{m}({\bf x})\Big|^{2}d{\bf x}+\int_{\partial\Omega}|\partial_{\nu}\varphi_{m}-\mathbf{f}_{m}|^{2}d\sigma({\bf x})+\epsilon\|\varphi_{m}\|_{H^{2}(\Omega)^{d}}^{2}\Bigg].$ (4.2)
@@
The existence and uniqueness of the minimizer of $J^{\epsilon,N}_{{\bf U}^{(k)}}$ are guaranteed by the strict convexity of the functional and the fact that its domain $H$ is a Hilbert space.
@@
We summarize the Picard iteration and quasi-reversibility method to solve (3.17) and hence Problem 1 in Algorithm 1.
@@
Algorithm 1 Solution procedure for Problem 1 via the time dimension reduction method
1:Select a cutoff number NN the regularization parameter $\epsilon$.
2:Set ${\bf U}^{(0)}={\bf 0}\in H.$
3:for $k=0$ to K1K-1, for some $K\in\mathbb{N}$ do
@@
4:  Minimize $J^{\epsilon,N}_{{\bf U}^{(k)}}$ defined in (4.2) and set ${\bf U}^{(k+1)}$ as the obtained minimizer.
5:end for
6:Set the computed solution ${\bf U}^{\rm comp}={\bf U}^{K}$:
$U^{\rm comp}=\begin{bmatrix}{\bf u}_{0}^{\rm comp}&{\bf u}_{1}^{\rm comp}&\dots&{\bf u}_{N}^{\rm comp}\end{bmatrix}^{\top}$.
@@
7:Reconstruct the fluid velocity field ${\bf u}^{\rm comp}$ using space-time expansion:
${\bf u}^{\rm comp}(\mathbf{x},t)=\mathbb{S}^{N}[U^{\rm comp}]=\sum_{n=0}^{N}{\bf u}_{n}^{\rm comp}(\mathbf{x})\Psi_{n}(t),$ (4.3)
for all $(\mathbf{x},t)\in\Omega\times(0,T)$.
8:Compute the approximate initial velocity field:
$\mathbf{U}_{0}^{\rm comp}(\mathbf{x})={\bf u}^{\rm comp}(\mathbf{x},0),$
for all $\mathbf{x}\in\Omega$.
@@
While our numerical experiments indicate that the iterative sequence $\{\mathbf{U}^{(k)}\}_{k\geq 0}$ converges reliably to the target solution when initialized from zero, we do not provide a theoretical justification of this convergence. The main difficulty lies in the lack of an appropriate Carleman estimate tailored to the anisotropic Navier–Stokes equations in the time-reduced framework. Developing such an estimate is highly nontrivial due to the complex structure of the fourth-order viscosity tensor. In the isotropic setting, where the viscosity tensor assumes a simpler form, one could potentially employ Carleman weight functions to construct and analyze the convergence of an analogous iterative scheme using existing Carleman estimates [9, 11]. However, since the focus of this work is on anisotropic media, we defer the theoretical convergence analysis for the isotropic case to future research.

### v4

@@ 3 The time-dimensional reduction method
For large $N$, due to Theorem 1, the spectral projection $\mathbb{P}^{N}$ approximatedly commutes with $\partial_{t}$ and with $({\bf u}\cdot\nabla){\bf u}$. In other words, “project-then-operate” and “operate-then-project” agree in the limit. This consideration motivates our choice of the exponentially weighted Legendre basis $\{\Psi_{n}\}_{n\geq 0}$. Other time bases (e.g., the exponentially weighted polynomial basis in $L^{2}(0,T)$; see [13, 28]) need not have this property and may therefore be unsuitable for deriving (3.10).
Corollary 1.
Assume that $\rho\in C^{1}(\overline{\Omega}\times[0,T])$ is known and that
${\bf u}\in H^{3}\bigl((0,T);H^{s}(\Omega)^{d}\bigr)$
for some integer $s>\frac{d}{2}+1$. Then
$\lim_{N\to\infty}\left\|\mathbb{P}^{N}[\rho({\bf x},t)\partial_{t}{\bf u}]-\rho({\bf x},t)\partial_{t}\mathbb{P}^{N}[{\bf u}]\right\|_{L^{2}((0,T);H^{s}(\Omega)^{d})}=0,$ (3.7)
and
$\lim_{N\to\infty}\left\|\mathbb{P}^{N}[\rho({\bf x},t)({\bf u}\cdot\nabla){\bf u}]-\rho({\bf x},t)(\mathbb{P}^{N}[{\bf u}]\cdot\nabla)\mathbb{P}^{N}[{\bf u}]\right\|_{L^{2}((0,T);H^{s-1}(\Omega)^{d})}=0.$ (3.8)
Proof.
To prove (3.7), we note first that
$\mathbb{P}^{N}[\rho({\bf x},t)\partial_{t}{\bf u}]\to\rho({\bf x},t)\partial_{t}{\bf u}\quad\text{in }L^{2}((0,T);H^{s}(\Omega)^{d}),$
by the approximation property of $\mathbb{P}^{N}$. On the other hand, by (3.4),
$\partial_{t}\mathbb{P}^{N}[{\bf u}]\to\partial_{t}{\bf u}\quad\text{in }L^{2}((0,T);H^{s}(\Omega)^{d}).$
Since $\rho({\bf x},t)$ is smooth, multiplication by $\rho({\bf x},t)$ is continuous on $H^{s}(\Omega)^{d}$. Therefore,
$\rho({\bf x},t)\partial_{t}\mathbb{P}^{N}[{\bf u}]\to\rho({\bf x},t)\partial_{t}{\bf u}\quad\text{in }L^{2}((0,T);H^{s}(\Omega)^{d}).$
Combining these two convergences, we obtain (3.7).
The proof of (3.8) is analogous. Indeed, to prove (3.8), we first note that
$\mathbb{P}^{N}[\rho({\bf x},t)({\bf u}\cdot\nabla){\bf u}]\to\rho({\bf x},t)({\bf u}\cdot\nabla){\bf u}\quad\text{in }L^{2}((0,T);H^{s-1}(\Omega)^{d}),$
by the approximation property of $\mathbb{P}^{N}$. On the other hand, by (3.5),
[…]
Again, since multiplication by the smooth coefficient $\rho({\bf x},t)$ is continuous, we obtain
$\rho({\bf x},t)(\mathbb{P}^{N}[{\bf u}]\cdot\nabla)\mathbb{P}^{N}[{\bf u}]\to\rho({\bf x},t)({\bf u}\cdot\nabla){\bf u}\quad\text{in }L^{2}((0,T);H^{s-1}(\Omega)^{d}).$
Hence (3.8) follows. ∎
Remark 6 (Regularity in Theorem 1, Corollary 1, and observed performance).
Theorem 1 and Corollary 1 require relatively high regularity of the solution. In particular, we assume
${\bf u}\in H^{3}\bigl((0,T);H^{s}(\Omega)^{d}\bigr)$
for some integer $s>\frac{d}{2}+1$, and in the corollary we additionally assume that the known density $\rho({\bf x},t)$ is sufficiently smooth. These assumptions are introduced to justify the projection-based analysis and to obtain clear convergence statements for both the unweighted and $\rho$-weighted terms.
In computations, however, the algorithm performs well even when the initial condition ${\bf u}^{0}$ belongs only to $L^{2}(\Omega)^{d}$ and may be discontinuous, so the assumptions of Theorem 1 and Corollary 1 are not satisfied; see Section 5 for our out-of-expectation numerical results. This suggests that the practical scope of the method may be substantially broader than what is currently justified by the available theory.
Due to Corollary 1 and the limits (3.7)–(3.8), the right-hand side of (3.3) vanishes asymptotically as $N\to\infty$. Therefore, equation (3.3) can be approximated by the reduced system
$\rho({\bf x},t)\partial_{t}\mathbb{P}^{N}[{\bf u}]-{\rm div}\left(\bm{\mu}:\nabla\mathbb{P}^{N}[{\bf u}]\right)+\rho({\bf x},t)(\mathbb{P}^{N}[{\bf u}]\cdot\nabla)\mathbb{P}^{N}[{\bf u}]\\ +\mathbb{P}^{N}[\nabla p({\bf x},t)]-\mathbb{P}^{N}[\rho({\bf x},t){\bf F}({\bf x},t)]=\mathbf{0},$ (3.9)
for $({\bf x},t)\in\Omega\times(0,T),$ which serves as the foundation of our time-dimensional reduction model.
For each $n\geq 0$, let
${\bf u}_{n}({\bf x})=\int_{0}^{T}e^{-2t}{\bf u}({\bf x},t)\Psi_{n}(t)\,dt,$ (3.10)
be the $n$-th Fourier mode of $\bf u$} with respect to the basis $\{\Psi_{n}\}_{n\geq 0}$.
@@
$\mathbb{P}^{N}[{\bf u}]({\bf x},t)=\sum_{n=0}^{N}{\bf u}_{n}({\bf x})\Psi_{n}(t).$ (3.11)
Substituting (3.11) into the system (3.9), we obtain the approximation
$\rho({\bf x},t)\sum_{n=0}^{N}{\bf u}_{n}({\bf x})\Psi_{n}^{\prime}(t)-\sum_{n=0}^{N}{\rm div}\left(\bm{\mu}:\nabla{\bf u}_{n}({\bf x})\right)\Psi_{n}(t)\\ +\rho({\bf x},t)\sum_{l=0}^{N}\sum_{n=0}^{N}\left({\bf u}_{l}({\bf x})\cdot\nabla\right){\bf u}_{n}({\bf x})\Psi_{l}(t)\Psi_{n}(t)+\mathbb{P}^{N}[\nabla p({\bf x},t)]-\mathbb{P}^{N}[\rho({\bf x},t){\bf F}({\bf x},t)]=\mathbf{0},$ (3.12)
valid for all $({\bf x},t)\in\Omega\times(0,T)$.
To derive the time-reduced model, we test equation (3.12) against $e^{-2t}\Psi_{m}(t)$ and integrate over $t\in(0,T)$ for each $m=0,1,\dots,N$. This yields
$\sum_{n=0}^{N}s_{mn}({\bf x}){\bf u}_{n}({\bf x})+\sum_{l=0}^{N}\sum_{n=0}^{N}a_{mnl}({\bf x})\left({\bf u}_{l}({\bf x})\cdot\nabla\right){\bf u}_{n}({\bf x})\\ =-\nabla p_{m}({\bf x})+{\rm div}\left(\bm{\mu}:\nabla{\bf u}_{m}({\bf x})\right)+{\bf F}_{m}({\bf x}),$ (3.13)
for all ${\bf x}\in\Omega$, where
$s_{mn}({\bf x})=\int_{0}^{T}e^{-2t}\rho({\bf x},t)\Psi_{n}^{\prime}(t)\Psi_{m}(t)\,dt,$
$a_{mnl}({\bf x})=\int_{0}^{T}e^{-2t}\rho({\bf x},t)\Psi_{l}(t)\Psi_{n}(t)\Psi_{m}(t)\,dt,$
@@
$p_{m}({\bf x})=\int_{0}^{T}e^{-2t}p({\bf x},t)\Psi_{m}(t)\,dt,\qquad{\bf F}_{m}({\bf x})=\int_{0}^{T}e^{-2t}\rho({\bf x},t){\bf F}({\bf x},t)\Psi_{m}(t)\,dt.$
The corresponding Cauchy boundary conditions for each mode ${\bf u}_{m}$ follow from (1.2) and (3.10):
${\bf u}_{m}({\bf x})=\mathbf{0},\qquad\partial_{\nu}{\bf u}_{m}({\bf x})=\int_{0}^{T}e^{-2t}\mathbf{f}({\bf x},t)\Psi_{m}(t)\,dt,$ (3.14)
for all ${\bf x}\in\partial\Omega$.
Equations (3.13) and (3.14) define a coupled boundary value problem for the vector of coefficients $\mathbb{P}^{N}[{\bf u}]$. We refer to this system as the time-dimensional reduction model:
$\begin{cases}\displaystyle\sum_{n=0}^{N}s_{mn}({\bf x}){\bf u}_{n}({\bf x})+\sum_{l=0}^{N}\sum_{n=0}^{N}a_{mnl}({\bf x})\left({\bf u}_{l}({\bf x})\cdot\nabla\right){\bf u}_{n}({\bf x})\\ \hskip 142.26378pt=-\nabla p_{m}({\bf x})+{\rm div}\left(\bm{\mu}:\nabla{\bf u}_{m}({\bf x})\right)+{\bf F}_{m}({\bf x}),&{\bf x}\in\Omega,\\[8.00003pt] \displaystyle{\bf u}_{m}({\bf x})=\mathbf{0},&{\bf x}\in\partial\Omega,\\[3.99994pt] \displaystyle\partial_{\nu}{\bf u}_{m}({\bf x})=\mathbf{f}_{m}({\bf x}):=\int_{0}^{T}e^{-2t}\mathbf{f}({\bf x},t)\Psi_{m}(t)\,dt,&{\bf x}\in\partial\Omega,\end{cases}$ (3.15)
@@
Solving the system (3.15) yields the coefficient functions ${\bf u}_{0},{\bf u}_{1},\dots,{\bf u}_{N}$, which can be used to reconstruct an approximation of the original space-time solution ${\bf u}({\bf x},t)$ via
$\mathbb{P}^{N}[{\bf u}]({\bf x},t)=\sum_{n=0}^{N}{\bf u}_{n}({\bf x})\Psi_{n}(t).$
According to the approximation formula (3.11), this reconstruction provides a finite-dimensional approximation of the original space-time solution $\bf u$}. Hence, the inverse problem under consideration (Problem 1) is transformed into the computational task of solving the reduced model (3.15).
Despite this reduction, solving the system (3.15) remains highly nontrivial due to the presence of the nonlinear convective term
$\sum_{l=0}^{N}\sum_{n=0}^{N}a_{mnl}({\bf x})\left({\bf u}_{l}({\bf x})\cdot\nabla\right){\bf u}_{n}({\bf x}),\qquad m\in\{0,1,\dots,N\},$ (3.16)
which makes (3.15) a coupled nonlinear system. To overcome this difficulty, we apply a conventional damped Picard iteration; see Section 4.
@@
The time-dimensional reduction model (3.15) is an approximate representation of the original system, arising from the truncation of the infinite Fourier expansion in (3.11). Nevertheless, we adopt the working assumption that
$\mathbf{U}=\begin{bmatrix}{\bf u}_{0}&{\bf u}_{1}&\dots&{\bf u}_{N}\end{bmatrix}^{\top}$
satisfies the reduced model. This assumption is justified by Theorem 1 and Corollary 1, which show that the projection and weighted commutation errors vanish as $N\to\infty$. Consequently, the reduced model becomes an increasingly accurate surrogate for the full dynamics as the cutoff parameter $N$ increases. From a computational viewpoint, this assumption is also natural, since any practical implementation necessarily uses a finite truncation level $N$.
@@
The uniqueness of the time-reduction model (3.15) remains an open problem due to the complexity introduced by the anisotropic fourth-order tensor $\bm{\mu}$ and the nonlinear convection term (3.16). A standard approach in inverse problems is to use Carleman estimates to establish uniqueness and stability. However, deriving such estimates in the fully anisotropic setting is highly nontrivial and, to the best of our knowledge, is not yet available in a general form. In contrast, in the isotropic case, Carleman estimates are more tractable; see, for instance, [9, 11]. Under such simplifications, the uniqueness of the time-reduction model (3.15) may become accessible to rigorous analysis. We leave this direction for future research.
4 A damped Picard algorithm to solve the time-reduction model
@@
which serves as the admissible space for the reduced system in (3.15). Let the initial guess be
@@
is known for some $k\geq 0$. We first compute an intermediate iterate
$\widetilde{\bf U}^{(k+1)}=\begin{bmatrix}\widetilde{{\bf u}}^{(k+1)}_{0}&\widetilde{{\bf u}}^{(k+1)}_{1}&\dots&\widetilde{{\bf u}}^{(k+1)}_{N}\end{bmatrix}^{\top}$
as the solution of the following linearized boundary value problem:
@@
$\begin{cases}\displaystyle\sum_{n=0}^{N}s_{mn}({\bf x})\widetilde{{\bf u}}_{n}^{(k+1)}({\bf x})-{\rm div}\left(\bm{\mu}:\nabla\widetilde{{\bf u}}_{m}^{(k+1)}({\bf x})\right)\\[3.99994pt] \displaystyle\hskip 56.9055pt=-\sum_{l=0}^{N}\sum_{n=0}^{N}a_{mnl}({\bf x})\left({\bf u}_{l}^{(k)}({\bf x})\cdot\nabla\right){\bf u}_{n}^{(k)}({\bf x})-\nabla p_{m}({\bf x})+{\bf F}_{m}({\bf x}),&{\bf x}\in\Omega,\\[8.00003pt] \widetilde{{\bf u}}_{m}^{(k+1)}({\bf x})=\mathbf{0},&{\bf x}\in\partial\Omega,\\[3.99994pt] \partial_{\nu}\widetilde{{\bf u}}_{m}^{(k+1)}({\bf x})=\mathbf{f}_{m}({\bf x}),&{\bf x}\in\partial\Omega,\end{cases}$ (4.1)
for each $m=0,1,\dots,N$.
We then define the damped Picard update by
${\bf U}^{(k+1)}=(1-\omega){\bf U}^{(k)}+\omega\,\widetilde{\bf U}^{(k+1)},\qquad\omega=0.5.$ (4.2)
This iteration is a damped fixed-point scheme, where the nonlinear convection term is evaluated at the previous iterate, and the damping parameter is used to improve stability.
@@
The next step is to solve the over-determined problem (4.1). By over-determined, we mean that both Dirichlet and Neumann boundary conditions are prescribed for $\widetilde{\bf U}^{(k+1)}$. We propose to use the quasi-reversibility method, first introduced in [18]. Thus, the intermediate iterate $\widetilde{\bf U}^{(k+1)}$ is defined as the minimizer of the least-squares functional
@@
$J_{{\bf U}^{(k)}}^{\epsilon,N}(\bm{\varphi})=\sum_{m=0}^{N}\Bigg[\int_{\Omega}\Bigg|{\rm div}\left(\bm{\mu}:\nabla\varphi_{m}({\bf x})\right)-\sum_{n=0}^{N}s_{mn}({\bf x})\varphi_{n}({\bf x})\\ -\sum_{l=0}^{N}\sum_{n=0}^{N}a_{mnl}({\bf x})\left({\bf u}_{l}^{(k)}({\bf x})\cdot\nabla\right){\bf u}_{n}^{(k)}({\bf x})-\nabla p_{m}({\bf x})+{\bf F}_{m}({\bf x})\Bigg|^{2}d{\bf x}\\ +\int_{\partial\Omega}\left|\partial_{\nu}\varphi_{m}({\bf x})-\mathbf{f}_{m}({\bf x})\right|^{2}d\sigma({\bf x})+\epsilon\|\varphi_{m}\|_{H^{2}(\Omega)^{d}}^{2}\Bigg].$ (4.3)
@@
The existence and uniqueness of the minimizer of $J_{{\bf U}^{(k)}}^{\epsilon,N}$ follow from the strict convexity of the functional and the fact that its domain $H$ is a Hilbert space.
@@
We summarize the damped Picard iteration and quasi-reversibility method for solving (3.15), and hence Problem 1, in Algorithm 1.
@@
Algorithm 1 Solution procedure for Problem 1 via the time-dimensional reduction method
1:Select a cutoff number NN and the regularization parameter $\epsilon$.
2:Set the damping parameter $\omega=0.5$.
3:Set ${\bf U}^{(0)}={\bf 0}\in H$.
4:for $k=0$ to K1K-1, for some $K\in\mathbb{N}$ do
@@
5:  Minimize the functional $J_{{\bf U}^{(k)}}^{\epsilon,N}$ defined in (4.3), and denote the obtained minimizer by $\widetilde{\bf U}^{(k+1)}$.
6:  Update the iterate by the damped Picard step
${\bf U}^{(k+1)}=(1-\omega){\bf U}^{(k)}+\omega\,\widetilde{\bf U}^{(k+1)}.$
7:end for
8:Set the computed solution ${\bf U}_{\rm comp}={\bf U}^{K}$:
${\bf U}_{\rm comp}=\begin{bmatrix}{\bf u}_{0,{\rm comp}}&{\bf u}_{1,{\rm comp}}&\dots&{\bf u}_{N,{\rm comp}}\end{bmatrix}^{\top}$.
@@
9:Reconstruct the fluid velocity field ${\bf u}_{\rm comp}$ using the space-time expansion
${\bf u}_{\rm comp}({\bf x},t)=\mathbb{S}^{N}[{\bf U}_{\rm comp}]=\sum_{n=0}^{N}{\bf u}_{n,{\rm comp}}({\bf x})\Psi_{n}(t),$ (4.4)
for all $({\bf x},t)\in\Omega\times(0,T)$.
10:Compute the approximate initial velocity field
${\bf u}^{0}_{\rm comp}({\bf x})={\bf u}_{\rm comp}({\bf x},0),\qquad{\bf x}\in\Omega.$
@@
While our numerical experiments indicate that the iterative sequence $\{{\bf U}^{(k)}\}_{k\geq 0}$ converges reliably to the target solution when initialized from zero, we do not provide a theoretical justification of this convergence. The main difficulty lies in the absence of an appropriate Carleman estimate tailored to the anisotropic Navier–Stokes equations in the time-reduced setting. Developing such an estimate is highly nontrivial due to the complex structure of the fourth-order viscosity tensor. In the isotropic case, where the viscosity tensor has a simpler form, one may hope to analyze a similar iterative scheme using existing Carleman estimates; see, for instance, [9, 11]. Since the present work focuses on anisotropic media, we leave the corresponding theoretical convergence analysis for future research.